OpenStudy (aivantettet26):

what is the derivative of sec^2y + cot^2x = 3?

OpenStudy (aivantettet26):

I got this right but i just got half-credit because the choices utilized identities

terenzreignz (terenzreignz):

Well let's see... show me what you did.

OpenStudy (aivantettet26):

sec^2(y)tan^2(y) -csc^2(x) = 0

OpenStudy (aivantettet26):

but from there on out. my prof labeled me wrong T.T

OpenStudy (aivantettet26):

the answer is cotxcsc^2x / sec^2ytany

terenzreignz (terenzreignz):

Well naturally :) Derivative of sec²(y) is 2sec²(y)tan(y) y'

terenzreignz (terenzreignz):

Let's differentiate \(\large \sec^2(y)\) using chain rules: \[\Large \implies \color{red}{2\sec(y)} \frac{d}{dx}\sec(y)\]\[\Large \implies 2\sec(y)\color{red}{\sec(y)\tan(y)}\frac{d}{dx}y\]\[\Large \implies 2\color{red}{y'}\sec^2(y)\tan(y)\] gets? ^_^

OpenStudy (aivantettet26):

oh look at that cheese! haha!

OpenStudy (aivantettet26):

where I really went wrong is with the identities

terenzreignz (terenzreignz):

You'd go wrong with them if you went wrong with differentiation in the first place. Now, did you understand the proper way to differentiate sec²(y) ?

OpenStudy (aivantettet26):

yeah. I messed up

terenzreignz (terenzreignz):

So we now have \[\Large 2\color{red}{y'}\sec^2(y)\tan(y )-2\cot(x)\csc^2(x)=0\]

terenzreignz (terenzreignz):

Do I have to go through the nitty gritty of differentiating cot²(x) ?

OpenStudy (aivantettet26):

no need i recognize my mistake

OpenStudy (aivantettet26):

wrong grammar pa haha

terenzreignz (terenzreignz):

hmm...

OpenStudy (aivantettet26):

2y′sec2(y)tan(y))=2cot(x)csc2(x)

OpenStudy (aivantettet26):

2y′=2cot(x)csc2(x)/sec2(y)tan(y))

OpenStudy (aivantettet26):

y'=cot(x)csc2(x)/sec2(y)tan(y)

terenzreignz (terenzreignz):

mhmm :) \[\Large \color{red}{y'}= \frac{\cot(x) \csc^2(x)}{\tan(y)\sec^2(y)}\]

OpenStudy (aivantettet26):

i fluttering hate myself for not getting this correct >.< rookie mistake

terenzreignz (terenzreignz):

Ahh don't worry about it ^_^

terenzreignz (terenzreignz):

We all make mistakes XD